Suppose that people who study very hard for the BAR Exam have an odds of 1.2 of passing the BAR Exam. Suppose that people who do not study for the BAR Exam have an odds of .34 of passing the BAR Exam. What is the Relative Risk for failing the BAR Exam comparing people who did not study to people who studied hard, round your answer to at least 2 decimal places?
Answer #1

The contingency table is,

Study No Study
Fail a b a+b
Pass c d c + d
a+c b+d

Given,

c/a = 1.2 => c/a + 1 = 1.2 + 1 => (a + c) / a = 2.2 => a/(a + c) = 1/2.2

d/b = 0.34 => d/b + 1 = 0.34 + 1 => (b + d)/ b = 1.34 => b / (b + d) = 1/1.34

Relative risk = (b / b+d) / (a/a + c)

= (1/1.34) / (1 / 2.2)

= 2.2 / 1.34

= 1.641791